OpenStudy (anonymous):

x/x-1=1/x+1+3/x^2-1 solve using appropriate technique????

OpenStudy (unklerhaukus):

\[\frac x{x-1}=\frac 1{x+1}+\frac{3}{x^2-1}\]

OpenStudy (unklerhaukus):

have i got this right ?

OpenStudy (unklerhaukus):

multiply by this \[(x-1)(x+1)=x^2-1\]

OpenStudy (anonymous):

Multiply by x^2-1 on both sides

OpenStudy (jiteshmeghwal9):